OpenStudy (anonymous):

Which factor can you cancel from the numerator and denominator to simplify 8-2x/3x-12? [Hint: (a−x) = − (x − a)]

OpenStudy (asnaseer):

can you notice a common factor in the numerator between the terms 8 and 2x?

OpenStudy (anonymous):

2

OpenStudy (asnaseer):

good, so can you write the numerator with the 2 taken out as a factor?

OpenStudy (anonymous):

sorry that kinda confused me

OpenStudy (asnaseer):

ok - you correctly found that 2 is a common factor between 8 and 2x, this means we can factor the numerator as follows:\[8-2x=2(4-x)\]agreed?

OpenStudy (anonymous):

oh yes. i agree

OpenStudy (asnaseer):

good - now try the same thing with the denominator - what is common between 3x and 12?

OpenStudy (anonymous):

3 so 2(x-4)

OpenStudy (anonymous):

*3

OpenStudy (asnaseer):

perfect! so now you end up with:\[\frac{8-2x}{3x-12}=\frac{2(4-x)}{3(x-4)}\]next, try and make use the hint given to you in the question.

OpenStudy (anonymous):

so i cold cancel out the 4's?

OpenStudy (asnaseer):

no, what you need to do is notice that:\[(4-x)=-(x-4)\]

OpenStudy (asnaseer):

do you understand that?

OpenStudy (anonymous):

not really

OpenStudy (asnaseer):

ok, let me try to explain...

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

\[-(x-4)=-x-(-4)=-x+4=4-x\]does it make sense now?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

ok, so if we use this fact, then we can substitute:\[(4-x)=-(x-4)\]into out fraction to get:\[\frac{8-2x}{3x-12}=\frac{2(4-x)}{3(x-4)}=\frac{-2(x-4)}{3(x-4)}\]

OpenStudy (asnaseer):

make sense?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

ok now you should be able to see that we can cancel out \((x-4)\) from numerator and denominator.

OpenStudy (anonymous):

makes sense now! thanks so much! tremendous help!

OpenStudy (asnaseer):

yw :)